# Shield Synthesis for LTL Modulo Theories

### v1

@@ 4.2 Temporal Dynamics
$\varphi_{0}$30.30.30.30 $631$$19$17.17.17.17 $171$ $173$ $183$ $184$
$\varphi_{1}$41.41.41.41 $590$22 .21 190 192 199 212
$\varphi_{2}$80.80.80.80 $520$24 .20 214 194 214 199
$\varphi_{3}$45.45.45.45 $340$22 .18 106 105 120 105
$\varphi_{4}$50.50.50.50 $640$21 .17 118 122 120 117
$\varphi_{5}$80.80.80.80 $520$21 .16 124 124 127 156
$\varphi_{6}$37.37.37.37 $370$16 .14 141 156 148 151
$\varphi_{7}$49.49.49.49 $710$20 .20 163 159 165 173
$\varphi_{8}$45.45.45.45 $580$21 .15 133 134 143 189
$\varphi_{9}$18.18.18.18 $610$14 .11 170 179 172 203
$\varphi_{10}$50.50.50.50 $690$25 .19 173 168 182 177
$\varphi_{11}$31.31.31.31 $530$21 .17 177 178 191 177
$\varphi_{12}$57.57.57.57 $510$29 .22 156 155 179 182
Table 2: Comparison of (?) and our approach (in $\mu s$)
We now show how $S_{\mathcal{T}}$ can manage both rich data and temporal dynamics. To do so, we used $13$ different specifications from (?), which are, to the best of our knowledge, the only previous successful attempt to compute rich-data shields. We generated our shield and measured time for computing: (i) the Boolean step (columns $\mathbb{B}$), and (ii) producing the final output (columns $\mathcal{T}$). We compared our results at Tab. 2. For the first task, our approach, on average, requires less than $0.21\mu{}s$, while (?) takes more than twice as long on average (with over $0.47\mu{}s$). Our approach was also more efficient in the second task, taking on average about $0.157ms$, whereas (?) required $0.557ms$ on average. Moreover, note that our ($\mathbb{B}$) contains both detection and correction of candidate outputs, whereas ($\mathbb{B}$) of  (?) only contains detection.
@@
It is important to note another key advantage of our approach:  (?) necessarily generates a WR, whereas we can also construct $C_{\mathcal{T}}$ (column $\mathbb{B}^{\prime}$), performs considerably faster (and can operate leveraging highly matured techniques, e.g., bounded synthesis). In addition to this, we (?) presents a monolithic method that to be used with specifications containing only linear real arithmetic $\mathcal{T}_{\mathbb{Z}}$. Hence, if we slightly modify $\mathcal{T}$ to $\mathcal{T}_{\mathbb{Z}}$, then the method in (?) cannot provide an appropriate shield, whereas we do (column $\mathcal{T}_{\mathbb{Z}}$) (neither for non-linear $\mathcal{T}$). In contrast, our method encodes any arbitrary $\exists^{*}\forall^{*}$ decidable fragment of $\mathcal{T}$ (e.g., non-linear arithmetic or the array property fragment (?)). Additionally, we can optimize the outputs of $S_{\mathcal{T}}$ with respect to different criteria, such as returning the smallest/greatest safe output (column $\mathcal{T}_{A}$) and the output closest to the candidate (column $\mathcal{T}_{B}$), for which we used Z3 with OMT (??). These experiments show the applicability of our approach.  [^4] Note that because of space limitations we could not include experiments over real robots in the main text (see suppl. material)
5 Related Work and Conclusion
Related Work.
Classic shielding approaches (???????) focus on properties expressed in Boolean LTL, and are incompatible for systems with richer-data domains: i.e., they need explicit manual discretization of the requirements. In the other hand, we have a sound procedure that directly takes $\textup{LTL}_{\mathcal{T}}$ specifications, which is fundamentally new. Other, more recent, work (???) proves that, under certain conditions, $\textup{LTL}_{\mathcal{T}}$ synthesis is decidable via abstraction methods, which can be understood as computing minterms to produce symbolic automata and transducers (?) from reactive specifications, and using antichain-based optimization, as suggested by (?). We complement and extend these techniques and demonstrate that they can act as a basis for shields synthesis modulo theories, which also contains an error detection phase whose design is essential for permissivity and which can optimize the output with respect to the erroneous candidate. Our work also includes an in-depth comparison to  (?), that present a competing method for synthesizing shields in the $\mathcal{T}$ of linear real arithmetic, which does not cover such an expressive fragment and only computes WR.
As for reactive synthesis, methods for infinite theories have also been proposed. Moreover, some of them consider a expressive temporal logics where not only multi-step properties can be defined, but also variables can convey information across-time  (?). However, they do not guarantee success (?) or termination (?????), and can be unsound for some cases (?) or may need guidance (?). Therefore, any shield synthesis techniques derived from these would be incomplete. Runtime enforcement (???) is an additional technique similar to shielding, but it is not compatible with reactive systems (?). In addition, supervisory control of discrete event systems (?) is based on invariant specifications and controllable events, and while reactive shield-synthesis additionally admits rich temporality, supervisory control cannot be applied to temporal settings (e.g., the one presented in Subsec. 4.2).
Other approaches for ensuring DNN safety rely on formal verification (?????????????????????) and Scenario-Based Programming (???). but they have limited scalability and expressivity.
@@ Conclusion.
In this work, we present the first general methods for shielding properties encoded in $\textup{LTL}_{\mathcal{T}}$. These allows engineers to guarantee the safe behavior of DRL agents in complex, reactive environments. Specifically, we demonstrate how shields can be computed from a controller and from a winning region. This is not simply a direct application of synthesis, instead, we can guarantee maximally and minimally permissive shields, as well as shields optimized with respect to objective functions in arbitrary decidable $\mathcal{T}$. We also empirically demonstrate its applicability.
A next step is to adapt shields modulo theories to probabilistic settings (e.g., shielding POMDPs) (????). Also, we plan on incorporating our shield along with complementary DNN verification engines. Another direction of future work relies on recent results for satisfiability of finite-trace $\textup{LTL}_{\mathcal{T}}$ (?), as we aim to identify fragments of $\textup{LTL}_{\mathcal{T}}$ with across-time data-transfer for which reactive synthesis is decidable. Future work may also include $\textup{LTL}_{\mathcal{T}}$ shields for multi-agent and planning scenarios (????), and combinations with uncertainty or continuous time (????). We want to adapt results to finite traces (?), since our agents either collide, reach the goal or have a time limit (unknown a priori).

### v2

@@ Introduction
$\varphi_{0}$ $30$ $631$ $19$ $17$ $171$ $173$ $183$ $184$
$\varphi_{1}$ $41$ $590$ 22 21 190 192 199 212
$\varphi_{2}$ $80$ $520$ 24 20 214 194 214 199
$\varphi_{3}$ $45$ $340$ 22 18 106 105 120 105
$\varphi_{4}$ $50$ $640$ 21 17 118 122 120 117
$\varphi_{5}$ $80$ $520$ 21 16 124 124 127 156
$\varphi_{6}$ $37$ $370$ 16 14 141 156 148 151
$\varphi_{7}$ $49$ $710$ 20 20 163 159 165 173
$\varphi_{8}$ $45$ $580$ 21 15 133 134 143 189
$\varphi_{9}$ $18$ $610$ 14 11 170 179 172 203
$\varphi_{10}$ $50$ $690$ 25 19 173 168 182 177
$\varphi_{11}$ $31$ $530$ 21 17 177 178 191 177
$\varphi_{12}$ $57$ $510$ 29 22 156 155 179 182
Table 1: Comparison of our approach with (Wu et al. 2019), measured in $0.1\mu s$ for $\mathbb{B}$ and in $\mu s$ for $\mathcal{T}$, $\mathcal{T}_{\mathbb{Z}}$, $\mathcal{T}_{A}$ and $\mathcal{T}_{B}$.
@@
The approach by (Wu et al. 2019) is, to the best of our knowledge, the only previous successful attempt to compute rich-data shields. We compared our work with theirs: for the $13$ different specifications from (Wu et al. 2019), we generated our shield and measured time for computing: (i) the Boolean step (clm. $\mathbb{B}$), and (ii) producing the final output (clm. $\mathcal{T}$). We compared our results at Tab. 1. For the first task, our approach, on average, requires less than $0.21\mu{}s$, while they take more than twice as long on average (with over $0.47\mu{}s$). Our approach was also more efficient in the second task, taking on average about $0.157ms$, whereas they take required $0.557ms$ on average. Moreover, note that our ($\mathbb{B}$) contains both detection and correction of candidate outputs, whereas their ($\mathbb{B}$) is only detection. It is important to note another key advantage of our approach:  (Wu et al. 2019) necessarily generates a WR, whereas we can also construct $C_{\mathcal{T}}$ (clm. $\mathbb{B}^{\prime}$), performs considerably faster (and can operate leveraging highly matured techniques, e.g., bounded synthesis). In addition to this, we (Wu et al. 2019) presents a monolithic method that to be used with specifications containing only linear real arithmetic $\mathcal{T}_{\mathbb{Z}}$. Hence, if we slightly modify $\mathcal{T}$ to $\mathcal{T}_{\mathbb{Z}}$, then the method in (Wu et al. 2019) cannot provide an appropriate shield, whereas we do (clm. $\mathcal{T}_{\mathbb{Z}}$). Indeed, our method encodes any arbitrary $\exists^{*}\forall^{*}$ decidable fragment of $\mathcal{T}$ (e.g., non-linear arithmetic or the array property fragment (Bradley, Manna, and Sipma 2006)). Additionally, we can optimize the outputs of $S_{\mathcal{T}}$ with respect to different criteria, such as returning the smallest/greatest safe output (clm. $\mathcal{T}_{A}$) and the output closest to the candidate (clm. $\mathcal{T}_{B}$). For this, we used Z3 with optimization (Bjørner, Phan, and Fleckenstein 2015). Tab. 1 shows that our approach can manage both rich data and temporal dynamics.
@@ Conclusion.
In this work, we present the first general methods for shielding properties encoded in $\textup{LTL}_{\mathcal{T}}$. These allows engineers to guarantee the safe behavior of DRL agents in complex, reactive environments. Specifically, we demonstrate how shields can be computed from a controller and from a winning region. This is not simply a direct application of synthesis, instead, we can guarantee maximally and minimally permissive shields, as well as shields optimized with respect to objective functions in arbitrary decidable $\mathcal{T}$. We also empirically demonstrate its applicability, (see (Corsi et al. 2024) for an actual application).
A next step is to adapt shields modulo theories to probabilistic settings (Pranger et al. 2021; Carr et al. 2023) and planning (Camacho, Bienvenu, and McIlraith 2019). We also want to leverage recent results on finite-trace $\textup{LTL}_{\mathcal{T}}$ (Geatti et al. 2023), to model more expressive shields.